what temp change will occur when 1200 grams of water absorbs 13,650 calories of energy


1
Expert's answer
2021-02-24T03:46:47-0500

Given,

Mass of water "m=1200g"

Energy "E=13650" calories"=13650\times 4.2=57330J"

Specific heat of water "c=4.18J\/gK"


Let "\Delta t" be the change in temprature


As we know, "Q=mc\Delta t"


"\Rightarrow 57330=1200\times4.18\times \Delta t\\\Rightarrow\Delta t =\dfrac{57330}{5016}"


"\Rightarrow \Delta t=11.42K"